# Remote Auditing: Design-Based Tests of Randomization, Selection, and Missingness with Broadly Accessible Satellite Imagery

### v1

@@ 1.2 Test Statistic
$T\;=\;\sum_{i\in\text{test}}\Big[A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\Big]\;-\;\sum_{i\in\text{test}}\Big[A_{i}\log q_{i}+(1-A_{i})\log(1-q_{i})\Big],$ (2)
@@ 1.3 Reference Distribution and $p$-Values
which is valid in finite samples for arbitrary test statistics provided the resampling respects $\Omega$ (Candes et al., 2018; Hennessy et al., 2016). This procedure is nonparametric (no asymptotics needed), transparently honors blocking/stratification, and accommodates any off-the-shelf learner as the statistic’s engine. See Appendix for a simple sketch of validity.
@@
An important implementation detail concerns re-fitting versus re-using the trained learner. For strict finite-sample validity, we must refit within each resample using the same cross-fitting protocol (assuming the learner depends on the labels). In moderate samples with simple learners, the computational burden is modest.
Because the remote audit compresses pre-treatment imagery into high-dimensional embeddings $\boldsymbol{\phi}_{i}\in\mathbb{R}^{d}$ (with $d$ often in the hundreds or thousands, arising from multispectral indices and off-the-shelf encoders), the primary computational burden is re-fitting learners within each resample while preserving the design-based calibration to $\Omega$. We can therefore precompute $\boldsymbol{\phi}$ once and apply variance-reduction devices that do not alter the reference distribution.
@@
The first variance reduction device is “common random numbers” (Wright and Ramsay Jr, 1979): fix the cross-fitting split $S$ and all model-training seeds across resamples so that variation in $T^{(b)}$ arises solely from the assignment draws $A^{(b)}\sim\Omega$. This keeps the statistic exchangeable with its resampled counterparts while achieving target Monte Carlo precision with far fewer iterations. Second, in designs that treat exactly half of the units within each block, pairing every draw with its blockwise complement implements an antithetic coupling that further reduces Monte Carlo variance, provided the complement map preserves $\Omega$. Finally, parallelization across $b$ is straightforward. These choices keep the high-dimensional, imagery-driven audit inexpensive and fast without diluting its guarantees: $T$ continues to measure how much the embeddings predict assignment, and its extremeness is still judged against the finite-sample distribution implied by the registered mechanism.
@@
What about multiple testing in the remote audit? If several embeddings or hyperparameters are considered (e.g., CLIP-like encoders, ViT, Swin; patch sizes; resolutions), remote audits should control the family-wise error rate (FWER) or the false discovery rate (FDR). A simple and powerful approach for FWER is the Westfall–Young max-$T$ correction (Westfall and Young, 1993): at each resampled assignment $A^{(b)}$, compute every model’s $T^{(b)}$ and record the maximum. Compare the observed $T$ for each model to this max distribution to obtain adjusted $p$-values. Alternatively, Bonferroni or Benjamini–Hochberg-style methods are available (Thissen, Steinberg and Kuang, 2002). We recommend preregistering the model set and correction rule.
Putting the statistic $T$ and its randomization-based reference under $\Omega$ together, power is likely greatest when deviations from the registered mechanism align with pre-treatment visual signals, when the number of units is larger, when the embedding–learner pair captures stable structure while being evaluated strictly out of sample, and when the design leaves genuine headroom for predictability (highly constrained blockwise quotas leave less to detect). These ingredients translate directly into implementation choices: preserve block totals or cluster labels when resampling under $\Omega$, and prevent spatial leakage in model evaluation by constructing folds that leave out spatial neighborhoods or entire clusters; in our case study we mirror the blocking structure in the folds. Clustered or geographically stratified designs therefore pose no challenge for validity—both the resampling and the cross-fitting simply inherit the design’s constraints.
When assignment depends on factors that satellites cannot see (e.g., patronage networks), the audit will have limited power and this should be reported plainly. By contrast, when assignment comoves with roads, settlement density, roof materials, or land cover—features that proxy accessibility and wealth and are reliably visible from space (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019; Burke et al., 2021)—the test is often informative even with moderate $n$. In practice, precompute $\boldsymbol{\phi}$ once, start with simple learners and spatially robust cross-fitting, and use the Westfall–Young max-$T$ correction to discipline exploration across representations and scales. A small $p$-value then reads as evidence that the realized assignment is atypical of $\Omega$; a large $p$-value indicates that, at the resolution of the imagery and models used, the audit detected no image-aligned departures from the registered design.
2 A Practical Workflow and Reporting Standards
@@
Having defined the statistic and its design-based calibration, we now turn to practice. The workflow below translates the conditional randomization test into a preregistrable recipe that can be run before, during, or after field mobilization.
@@
Record the experiment’s randomization mechanism $\Omega$: complete/stratified/clustered, treatment fractions, and any constraints. If stratified, list strata membership for each unit. These inputs define the resampling.
@@
Select images that unambiguously precede any treatment or mobilization. When archives are sparse or cloudy, use compositing or median mosaics across pre-treatment windows. Avoid sensors whose earliest availability is post-treatment. Using post-treatment imagery risks mediator/collider bias if later repurposed in outcome models (Angrist and Pischke, 2009; Pearl, 2009). Archive scene IDs and acquisition dates in the replication package.
Step 3: Construct embeddings.
Choose interpretable indices (e.g., vegetation, built-up proxies) and/or general-purpose backbones (CLIP-like encoders, ViT, Swin) that have been used in Earth observation (Li et al., 2020; Dosovitskiy et al., 2020; Liu et al., 2021). Pretraining on remote-sensing corpora can help but is not required. Document patch size $s$, resolution, bands, and normalizations. Because embeddings are generated variables, keep them fixed throughout the audit and share code for reproducibility (Battaglia et al., 2024).
Step 4: Fit the predictive model with sample-splitting.
Use simple learners first (e.g., tree-based models), then consider neural models. Always evaluate out-of-sample (held-out fold or cross-fitting). Save the likelihood-based statistic $T$.
Step 5: Resample under $\Omega$ and compute the max-$T$.
Draw $B$ assignment vectors consistent with $\Omega$ (e.g., B=1, [^0] B=1{,}000), recompute $T^{(b)}$. If multiple image embedding representations are used, record the maximum test statistic across model per resample. Report adjusted $p$-values and show the observed $T$ against the reference distribution.
Step 6: Interpret cautiously and report transparently.
@@
Auxiliary audits.
The same logic supports audits of (i) selection into the experiment (predicting membership among a broader frame) and (ii) missingness (predicting which units have missing variables). These are not design-based in the same sense as the randomization audit because the resampling reference is less well pinned down; we therefore treat them as descriptive early-warning diagnostics that can motivate reweighting, oversampling, or targeted field follow-up.
@@
$\square$ Ethics & transparency Avoid post-treatment data; ensure geospatial privacy and responsible disclosure.
@@
Selection into the study frame.
Let $S_{i}\in\{0,1\}$ indicate whether unit $i$ in a broader, policy-relevant universe is enrolled in the experimental sample. When $\boldsymbol{\phi}_{i}$ strongly predicts $S_{i}$, the enrolled sample differs systematically from the target universe along pre-treatment features that are visible from space, raising an external-validity warning even if within-sample randomization is sound. We operationalize this as a covariate-shift diagnostic: train a classifier to distinguish enrolled units from units drawn from the putative frame using only $\boldsymbol{\phi}$, evaluate strictly out of sample, and summarize fit via likelihood improvement relative to the marginal sampling rate, and permute or randomly re-draw the $S_{i}$ selection indicator. Because there is no registered or otherwise defensible resampling mechanism for $S$ analogous to $\Omega$, these quantities are reported as descriptive diagnostics rather than design-valid tests.
Missingness and data quality.
When analysis requires adjustment, the same response models can be pre-specified as building blocks for principled corrections that do not rely on post-treatment information. For variables where missingness is plausibly at random given $\boldsymbol{\phi}$, inverse-probability weights $w_{ij}=R_{ij}/\hat{\rho}_{ij}(\boldsymbol{\phi}_{i})$ or multiple imputation models that condition on $\boldsymbol{\phi}$ provide transparent remedies; doubly robust procedures that combine a response model with an outcome model can be declared in the pre-analysis plan and implemented without altering the design-based logic of the randomization audit (Seaman and Vansteelandt, 2018). Where missingness is likely non-ignorable even after conditioning on $\boldsymbol{\phi}$, the imagery-based diagnostics still add value by localizing the problem and motivating sensitivity analyses and targeted re-contact. As with selection, we report these quantities as diagnostics rather than as hypothesis tests, and we apply the same cross-validation, sample-splitting, and multiplicity discipline used elsewhere in the audit.
In sum, selection and missingness audits repurpose the remote-audit machinery to answer two practical questions that matter for interpretation and planning: who entered the study relative to the intended universe, and where the data are likely to be thin. Their outputs—likelihood improvements, calibrated risk scores, overlap assessments, and, when warranted, transport or response weights—are easily preregistered, cheap to compute, and actionable during fieldwork. They do not substitute for theory or for direct measurement, and they are not design-valid tests in the sense of the conditional randomization procedure. They are, instead, early-warning indicators that help align designs with their targets and protect inference against predictable patterns of data loss.
3 Case Study: A Remote Audit of the Youth Opportunities Program (Uganda)
@@ Results.
Figure 3 reports three diagnostics. The randomization audit (Panel A) shows the observed likelihood-improvement statistic (vertical line) within the reference distribution under $\Omega$; the $p$-value is greater than 0.05, consistent with random assignment not being more predictable from pre-treatment imagery than chance. Two auxiliary checks are also informative. The selection audit (Panel B) contrasts YOP villages against a national frame: image features sharply distinguish enrolled vs. unenrolled locations, suggesting external validity concerns if these differences moderate treatment effects. Finally, the missingness audit (Panel C) indicates that pre-treatment features predict which units have missing variables, suggesting non-random data loss that merits attention in analysis plans (e.g., pre-specified handling of attrition).
@@
In this application, the remote audit would have (i) corroborated the lottery before expensive baseline enumeration, (ii) offered an early warning about representativeness (selection audit), and (iii) prompted pre-analysis plans for handling missing data differentially by location. None of these requires survey covariates. The audit also does not preclude standard balance tests once baselines are collected. Instead, it provides a low-cost, early-stage diagnostic that can be embedded in preregistration and updated as imagery and geocodes are refined.
4 Case Study: Detecting Faulty Randomization Based on a Retracted RCT in Bangladesh
We now apply the remote audit to Begum, Grossman and Islam (2022). Bonander et al. (2025) have raised substantial concerns about the randomization in this setting—including evidence that treatment assignment coincided perfectly with administrative boundaries and that the randomization is identical to some of the study authors used in prior, now retracted work (Islam, 2019). Our audit asks a narrower, design-based question: is the realized assignment in Begum, Grossman and Islam (2022) more predictable from pre-existing local conditions visible from satellite imagery than draws from the study’s purported randomization? The answer is yes, and strongly so, which is consistent with the broader record of irregularities summarized in Bonander et al. (2025). If the answer is yes, then investigators could have deployed the remote audit early on to identify problems with the RCT implementation before incurring the expense of on-the-ground data collection.
@@
The CRT does not identify who or what induced the deviation, nor does it imply that imagery features were used in implementation. It establishes a finite-sample discrepancy: under the claimed design, assignment should not be recoverable from pre-existing landscape signals; yet it is. Coupled with the independent evidence of geographically clustered assignment and shared treatment vectors across linked projects, the audit supports the conclusion that the realized allocation in Begum, Grossman and Islam (2022) was inconsistent with the stated randomization protocol. As with any imagery-only diagnostic, undetected deviations that are orthogonal to what satellites can see remain possible; here, however, two distinct learners point to the same rejection, and the effect sizes are large relative to the null. This analysis shows how aspects of randomization integrity can be tested before any fieldwork begins using the remote audit, successfully detecting possible irregularities noted in the replication community.
@@
5 What Else Should Analysts Check in Remote Audits?
Design-valid $p$-values are a starting point rather than a stopping rule. Because a remote audit inevitably reflects a handful of investigator choices, it is natural to ask whether the signal persists under reasonable alternatives. The brief checks below are quick to run and help separate substantive evidence from artifacts of modeling or preprocessing.
Vary the statistic and the data split.
Investigators can evaluate alternative strictly proper scores—such as Brier improvement or average cross-entropy per unit—to confirm that conclusions do not hinge on the score’s scale. It is also prudent to examine stability across random seeds and cross-fitting schemes, including spatially robust folds (e.g., leave-one-region-out or leave-one-cluster-out). Reporting the dispersion of $T$ (for example, an interquartile range across repeats) makes the degree of stability transparent.
Probe scale, sensors, and representation.
Choices about spatial scale and sensing are often consequential. Analysts may explore a range of patch sizes $s$ that plausibly cover geocoding error and local spillovers, and, where multiple pre-treatment sensors are available, run parallel analyses for each (with bandpass harmonization as needed). Comparing interpretable indices (vegetation, built-up, texture) with learned encoders helps distinguish model idiosyncrasies from image-aligned signal and can ease communication with non-technical audiences when the two agree.
Placebos and stress tests.

### v2

@@ 2.2 Test Statistic: Out-of-Sample Likelihood Improvement
$(T=)\;\Delta\mathcal{L}^{\star}\;\equiv\;\underbrace{\sum_{i\in\text{test}}\Big[A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\Big]}_{\mathcal{L}}\;-\;\underbrace{\sum_{i\in\text{test}}\Big[A_{i}\log q_{i}+(1-A_{i})\log(1-q_{i})\Big]}_{\mathcal{L}_{0}^{*}}.$ (2)
The “no-signal” anchor is here again 0; in finite samples, values below 0 are, as just discussed, not uncommon. The same construction underlies the auxiliary audits, swapping the treatment with missingness and selection into study indicators.
@@ 2.3 Reference Distribution and $p$-Values
which is valid in finite samples for arbitrary test statistics provided the resampling respects $\Omega$ (Candes et al., 2018; Hennessy et al., 2016). This procedure is nonparametric, transparently honors blocking/stratification, and accommodates any off-the-shelf learner as the engine of the statistic. See Appendix for a simple sketch of validity.
@@
An important implementation detail concerns re-fitting versus re-using the trained learner while remote auditing. For strict finite-sample validity, we must refit within each resample using the same cross-fitting protocol (assuming the learner depends on the labels). In moderate samples with simple learners and simple image features, such as NDVI (a measure of vegetation), the computational burden is modest.
@@
That said, when the remote audit compresses pre-treatment imagery into high-dimensional embeddings $\boldsymbol{\phi}_{i}\in\mathbb{R}^{d}$ (with $d$ often in the hundreds or thousands, arising from off-the-shelf encoders), the primary computational burden is re-fitting learners within each resample while preserving the design-based calibration to $\Omega$. We can therefore precompute $\boldsymbol{\phi}$ once and apply variance-reduction devices that do not alter the reference distribution. For example, one variance reduction device that can reduce the number of Monte Carlo iterations needed is “common random numbers” (Wright and Ramsay Jr, 1979): fix the cross-fitting split and all model-training seeds across resamples so that variation in $T^{(b)}$ arises solely from the assignment draws $A^{(b)}\sim\Omega$. This keeps the statistic exchangeable with its resampled counterparts while achieving target Monte Carlo precision with fewer iterations. Second, in designs that treat exactly half of the units within each block, pairing every draw with its blockwise complement implements an antithetic coupling that further reduces Monte Carlo variance, provided the complement map preserves $\Omega$. Finally, parallelization across $b$ is straightforward. These choices keep the high-dimensional, imagery-driven audit inexpensive and fast without diluting its guarantees.
@@
What about multiple testing in the remote audit? If several embeddings or hyperparameters are considered (e.g., CLIP-like encoders, ViT, Swin; patch sizes; resolutions), remote audits should control the family-wise error rate (FWER) or the false discovery rate (FDR). A simple and powerful approach for FWER is the Westfall–Young max-$T$ correction (Westfall and Young, 1993): at each resampled assignment $A^{(b)}$, compute every model’s $T^{(b)}$ and record the maximum. Compare the observed $T$ for each model to this max distribution to obtain adjusted $p$-values. Alternatively, Bonferroni or Benjamini–Hochberg-style methods are available (Thissen, Steinberg and Kuang, 2002). We recommend preregistering the model set and correction rule if used.
Putting the test statistic and its reference distribution together, power will be higher when deviations from the registered mechanism align with pre-treatment visual signals and when the embedding/learner captures stable structure under out-of-sample evaluation. If assignment turns on factors that satellites cannot see (e.g., patronage networks), the audit will have limited power. By contrast, when assignment covaries with roads, settlement density, roof materials, or land cover—features that proxy accessibility and wealth and are reliably visible from space (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019; Burke et al., 2021)—the test will be more informative.
Auxiliary audits.
As noted, similar logic supports audits of (i) selection into the experiment (predicting membership among a broader frame) and (ii) missingness (predicting which units have missing variables). These are not design-based in the same sense as the randomization audit because the resampling reference is less well pinned down; we therefore treat them as descriptive early-warning diagnostics that can motivate reweighting, oversampling, or field follow-up.
Selection into the study frame.
Let $S_{i}\in\{0,1\}$ indicate whether unit $i$ in a broader, policy-relevant universe is enrolled in the experimental sample, or not. When $\boldsymbol{\phi}_{i}$ strongly predicts selection into the study, $S_{i}$, the enrolled sample differs systematically from the target universe along pre-treatment features that are visible from space, raising an external-validity warning even if within-sample randomization is sound. We operationalize this as a covariate-shift diagnostic: train a classifier to distinguish enrolled units from units drawn from the putative frame using only $\boldsymbol{\phi}$, evaluate strictly out of sample, and summarize fit via likelihood improvement relative to the marginal sampling rate, and permute or randomly re-draw the $S_{i}$ selection indicator. Because there is no registered or otherwise defensible resampling mechanism for $S_{i}$ analogous to $\Omega$, these quantities are reported as descriptive diagnostics rather than design-valid tests.
Missingness and data quality.
When analysis requires adjustment, the same response models can be pre-specified as building blocks for principled corrections that do not rely on post-treatment information. For variables where missingness is plausibly at random given $\boldsymbol{\phi}$, inverse-probability weights $w_{ij}=R_{ij}/\hat{\rho}_{ij}(\boldsymbol{\phi}_{i})$ or multiple imputation models that condition on $\boldsymbol{\phi}$ provide transparent remedies (Blackwell, Honaker and King, 2017; Honaker, King and Blackwell, 2011); doubly robust procedures that combine a response model with an outcome model can be declared in the pre-analysis plan and implemented without altering the design-based logic of the randomization audit (Seaman and Vansteelandt, 2018). Where missingness is likely non-ignorable even after conditioning on $\boldsymbol{\phi}$, the imagery-based diagnostics still add value by localizing the problem and motivating sensitivity analyses and targeted re-contact. As with selection, we report these quantities as diagnostics rather than as hypothesis tests, and we apply the same cross-validation, sample-splitting, and multiple-testing discipline used elsewhere in the audit.
Randomization audit Selection audit Missingness audit
What it probes Realized assignment vs. declared design; pre-treatment covariates; blocks/clusters Who/what entered the study vs. target population; covariates and inclusion indicators Observed vs. missing outcomes/units; covariates and missingness indicator
Key assumption (focus) Under the registered mechanism, assignment is independent of pre-treatment features (within design) Inclusion unrelated to pre-treatment features after stated recruitment rules (diagnostic) Missingness unrelated to pre-treatment features (diagnostic)
Inferential goal Design-valid check: Is assignment unusually predictable vs. the randomization reference? Descriptive diagnostic: Is inclusion systematically predictable from features? Descriptive diagnostic: Is missingness systematically predictable from features?
Typical statistic Predict $A_{i}$ from $\mathbf{M}_{i}$; compare to permutation/CRT under $\Omega$ Predict inclusion $S_{i}$ from $\mathbf{M}_{i}$; quantify predictive strength/stability Predict missingness $R_{ij}$ from $\mathbf{M}_{i}$; quantify predictive strength/stability
Primary threat Deviations/manipulations of randomization; hidden stratifications Selection bias from sampling/consent/frame Attrition/nonresponse distorting the analysis set
When to use Pre/post implementation to verify assignment integrity During sampling/recruitment; external/internal selection concerns During collection/cleaning; nontrivial attrition/nonresponse
Actionable follow-ups Re-randomize/re-block; document deviations; sensitivity Reweight/adjust recruitment; bounds; document frame Weighting/imputation; bounds; follow-up for outcomes
Table 1: Contrasting randomization, selection, and missingness audits.
3 A Preregisterable Workflow
@@
Having defined the statistic and its design-based calibration, we now turn to practice. The workflow below translates the conditional randomization test into a preregistrable recipe that can be run before, during, or after field mobilization; mutatis mutandis, the same steps—swapping the label from assignment to sample membership or response status—produce the selection and missingness diagnostics reported later.
@@
Record the experiment’s randomization mechanism $\Omega$: treatment fractions, complete, stratified, or clustered structure, and other constraints. If stratified, list strata membership for each unit. These inputs define the resampling.
@@
Select images that unambiguously precede any treatment or mobilization. When archives are sparse or cloudy, use compositing or median mosaics across pre-treatment windows. Avoid sensors whose earliest availability is post-treatment for retrospective analyses (e.g., high-resolution Sentinel 2 data from European Space Agency becomes available only in 2015). Using post-treatment imagery risks mediator/collider bias if later repurposed in outcome models (Angrist and Pischke, 2009; Pearl, 2009; Cinelli, Forney and Pearl, 2024). Archive scene IDs and acquisition dates in the replication package. Clearly define patch size, resolution, bands, normalizations, and features used in processing.
Step 3: Fit the predictive model with sample-splitting.
Use simple learners first (e.g., tree-based models) if the sample size is small; otherwise, consider neural models using satellite imagery or image-derived features to predict treatment. Evaluate out-of-sample (held-out fold or cross-fitting). Save the likelihood-based statistic $T$.
Step 4: Resample under $\Omega$ and compute the max-$T$.
Draw $B$ assignment vectors consistent with $\Omega$ (e.g., B=1, [^0] B=1{,}000), recompute $T^{(b)}$. If multiple image embedding representations are used, record the maximum test statistic across models per resample. Report adjusted $p$-values and show the observed $T$ against the reference distribution.
Step 5: Interpret cautiously and report transparently.
@@
$\square$ Ethics & transparency Ensure anonymity in replication package as necessitated by ethical standards and required by IRB protocols.
@@
4 Case Study 1: A Remote Audit of the Youth Opportunities Program in Uganda
@@ Results.
Figure 3 reports three diagnostics. The randomization audit (Panel A) shows the observed likelihood-improvement statistic (vertical line) within the reference distribution under $\Omega$; the $p$-value is greater than 0.05, consistent with random assignment not being more predictable from pre-treatment imagery than chance. The two auxiliary checks are here informative. The selection audit (Panel B) contrasts YOP villages against a national frame: image features sharply distinguish enrolled vs. unenrolled locations, suggesting external validity concerns if these differences moderate treatment effects (Findley, Kikuta and Denly, 2021). Finally, the missingness audit (Panel C) indicates that pre-treatment features predict which units have missing variables, suggesting non-random data loss that merits attention in analysis plans (e.g., pre-specified handling of attrition).
@@
In this application, the remote audit would have (i) corroborated the lottery before expensive baseline enumeration, (ii) offered an early warning about representativeness (selection audit), and (iii) prompted pre-analysis plans for handling missing data differentially by location. None of these conclusions requires survey-collected covariate features. Of course, this audit also does not preclude standard balance tests once baseline data are, in fact, collected; rather, it provides a fast and low-cost early-stage diagnostic that can be embedded in preregistration (Lupia and Elman, 2014).
5 Case Study 2: Detecting Possibly Faulty Randomization Based on a Retracted RCT in Bangladesh
We next apply the remote audit to Begum, Grossman and Islam (2022). In this case, the central scientific question is whether the realized assignment conforms to the study’s registered mechanism $\Omega$. We restrict attention here to the design‑valid audit of randomization because independent work by Bonander et al. (2025) has raised concerns about this setting—including evidence that treatment assignment coincided with administrative boundaries and that the assignment vector is identical to one used by some study authors in prior, now‑retracted work (Islam, 2019). Our conditional randomization test asks a narrower question that fits our framework: is the realized assignment in more predictable from strictly pre‑treatment local conditions visible in satellite imagery than would be expected under draws from the purported design $\Omega$? The answer is yes, and strongly so—evidence that the observed allocation is atypical of the stated mechanism and consistent with the independent irregularities summarized by Bonander et al. (2025). Because the audit uses only pre‑treatment imagery and the registered design, it could have been deployed ex ante to flag problems with the RCT’s implementation before costly on‑the‑ground data collection.
@@
The CRT does not identify who or what induced the deviation, nor does it imply that imagery features were used in implementation. It establishes a finite-sample discrepancy: under the claimed design, assignment should not be recoverable from pre-existing landscape signals; yet it is. Coupled with the independent evidence of geographically clustered assignment and shared treatment vectors across linked projects, the audit is consistent with the conclusion that the realized allocation in Begum, Grossman and Islam (2022) may have deviated from the stated randomization protocol. As with any imagery-only diagnostic, further deviations that are unrelated to what satellites can see remain possible. To conclude, this analysis shows how aspects of randomization integrity can be tested before any fieldwork begins using the remote audit, detecting possible irregularities noted in the replication community.
@@
6 Discussion
Because any remote audit reflects a small number of investigator choices, analysts may naturally ask whether the signal persists under reasonable alternatives. Two quick and informative checks are to vary the strictly proper score (e.g., compare likelihood improvement to Brier improvement) and to assess stability across random seeds and cross-fitting schemes, including spatially robust folds such as leave-one-region-out or leave-one-cluster-out. Reporting the dispersion of the summary statistic $(T)$ across repeats makes the degree of stability transparent.